If a galaxy is 165 Mly away from us, how fast do we expect it to be moving and in what direction?

Answer in km/s
Answer #1

Here ,

distance , D = 165 Mly

Now , speed of galaxy = Ho*D

speed of galaxy = 20 * 165

speed of galaxy = 3300 km/s

speed of galaxy = 3.3 *10^4 km/s away from us